TAMENESS AND GEODESIC CORES OF SUBGROUPS

RITA GITIK

Abstract. We give a short new proof of the fact that quasiconvex subgroups of negatively curved groups are tame. The proof utilizes the concept of the geodesic core of the subgroup and is related to the Dehn algorithm.